the length of a rectangle is 3 less than twice the width. if the perimeter is 72cm, what is the lenght of the rectangle? im trying to help my son with his homework and dont know how to explain this
Answer by ankor@dixie-net.com(22740)

the length of a rectangle is 3 less than twice the width. if the perimeter is 72cm, what is the lenght of the rectangle?
:
Write an equation for each statement:
"the length of a rectangle is 3 less than twice the width. "
L = 2W - 3
:
"the perimeter is 72cm,"
2L + 2W = 72
This eq can be simplified, divide thru by 2
L + W = 36
:
From the 1st statement, replace L with (2W-3) in the above equation
(2W-3) + W = 36
add 3 to both sides
2W - 3 + 3 + W = 36 + 3
3W = 39
divide both sides by 3
W = 13 is the width
then
L = 2W - 3
L = 2(13) - 3
L = 26 - 3
L = 23 is length
:
:
Check this in the perimeter equation
2(23) + 2(13) =
46 + 26 = 72